The area of a rectangle is 91 square centimeters. Its length is x+2 and its width is x-4. Find the perimeter of the rectangle.
Found 2 solutions by Aldorozos, addingup:
Answer by Aldorozos(172)

(x+2)(x-4)= 90
x^2 -2x-8=90
solving this quadratic equation we get x=11
Width = 11-4 = 9
Length = 11+2 = 13
Perimeter is 2(9+13) = 2*22 = 44

Answer by addingup(3677)

L*W = 91
L = x+2
W = x-4
(x+2)*(x-4) = 91
x^2-2x-8 = 91
x^2-2x = 99
x^2-2x+1 = 100
(x-1)^2 = 100
x-1 = 10 or x-1 = -10
x = 11 or x = -9
Toss out the -9, we can't use it, and you have:
x = 11
The length is 11+2 = 13
The width is 11-4 = 7
The perimeter is:
2(13)+2(7)= 40
`````````````````````````````````````````
Check:
13+7 = 91 Correct